OpenStudy (anonymous):

Exactly 2/5 of the university's students live off campus. Of those students 3/10 of them live at home. What fraction of the university's students live at home?

OpenStudy (campbell_st):

well its 2/5 x 3/10 =

OpenStudy (anonymous):

\[\frac{2}{5} \times \frac{3}{10}=\frac{6}{50}=\frac{3}{25}=0.12\]

OpenStudy (anonymous):

Thank you